determine the magnitude of a magnetic force produced on a straight wire 60.0 cm long when a 10.5 A current flow on it while being exposed on a 27.5 T magnetic field acting 90.0^(0) on the wire.


1
Expert's answer
2022-05-23T10:17:54-0400
"F=BIl\sin\theta\\
=27.5*10.5*0.600*\sin90.0^\circ=173\:\rm N"